Can someone please help solve the following system of equations
2y=x+6
y=5+4x-x^2

Answer by mananth(16949)

2y=x+6..........1
y=5+4x-x^2 .....2
substitute y in (1)
2*(5+4x-x^2)=x+6
10+8x-2x^2=x+6
-2x^2+7x+4=0 (divide by -1)
2x^2-7x-4=0
2x^2-8x+x-4=0
2x(x-4)+1(x-4)=0
(2x+1)(x-4)=0
x=4 OR x=-1/2
Plug the values of x in (1) 2y=x+6
2y=10 so x=5
(4,5)
2y=x+6 x=-1/2
2y=-1/2+6
2y=11/2
y=11/4
(-1/2,11/4)
m.ananth@hotmail.ca